# Ground Every Sentence: Improving Retrieval-Augmented LLMs with Interleaved Reference-Claim Generation

### v1

@@ Methods
EM Rec. Exact match recall rate of gold short answers in the text generated by the model.
Claim Rec. Recall rate of generated claims in the text generated by the model.
@@
Table 5: A notation table.
In Table 5, we list the notations and abbreviations in this paper, along with their definitions.
@@ Appendix B Details of Training Dataset Construction
B.1 Citation Selection
During the training data structuring phase, we initially employ ChatGPT to automatically segment the answers within the WebGLM-QA dataset into clauses. Subsequently, for each of these clauses, we undertake a search for corresponding sentence-level, fine-grained citations from the provided reference passages. The prompt we use is as follows:
@@ B.2 Citation Filtering
Appendix C Experiment Settings
C.1 Details of Datasets
The statistic of the test datasets of ReClaimis shown in Table 7.
ASQA
ASQA Stelmakh et al. (2022) is the first long-form question answering dataset that focuses on ambiguous factoid questions (see Table 9). It contains 4,353 samples for the train split and 948 samples for the dev split. For ALCE Gao et al. (2023b), ASQA questions were paired with passages from Wikipedia passages (2018-12-20 snapshot) which purportedly contained the answers.
ELI5
ELI5 Fan et al. (2019) dataset is a benchmark in natural language processing designed for long-form question answering tasks, focusing on complex and explanatory questions (see Table 9). It comprises 270K threads from the Reddit forum “Explain Like I’m Five” (ELI5) where an online community provides answers to questions which are comprehensible by five year olds. For ALCE Gao et al. (2023b), ELI5 questions were paired with passages from Sphere Piktus et al. (2021), a filtered version of Common Crawl.
@@ C.1 Details of Datasets
ASQA 948 Factoid (ambiguous) Wikipedia (21M) 9.0 500.0 71.8
ELi5 1000 Why/How/What Sphere (899M) 16.5 500.0 121.5
@@
ASQA Question: Who has the highest goals in world football?
Reference Passages: [1] "Football Player of the Century", by IFFHS International Federation of Football History and Statistics, 1999, "South America Football Player of the Century", by IFFHS International Federation of Football History and Statistics. Pelé’s 1281 goals are recognized by FIFA as the highest total achieved by a professional footballer, although the Soccer Statistic Foundation (rssf) recognizes only 767 goals in official mode, occupying the third place after Josef Bican (805) and Romario (772). For his part, Maradona has been named the best soccer player in World Cup history both by The Times and FourFourTwo, publication that also rewarded him as the "Best
[2] for Christmas, but died less than 2 weeks before that, at the age of 88. Josef Bican Josef "Pepi" Bican (25 September 1913 —— 12 December 2001) was a Czech-Austrian professional footballer who played as a striker. Rec.Sport.Soccer Statistics Foundation (RSSSF) estimates that he scored at least 805 goals in all competitive matches, which would make him the most prolific scorer of all time. Having scored a record 600 league goals and at least 1468 goals overall, the International Federation of Football History & Statistics (IFFHS) awarded Bican the "Golden Ball" as the greatest goalscorer of the last century. He
[3] …
Answer: The players with the highest all-time goals and highest men’s and women’s international football goals differ. The player with the highest all-time men’s football goals is Josef Bican, who in 2020 was recognized by FIFA, the international governing body of football, as the record scorer with an estimated 805 goals. Christine Sinclair has the highest goals in women’s international football with 187 and is the all-time leader for international goals scored for men or women. Cristiano Ronaldo and Ali Daei are currently tied for leading goalscorer in the history of men’s international football with 109.
ELI5 Question: How are firms like snapchat, uber etc valued so highly while still not making a profit? Do venture capitalists not expect some form of repayment within a number of years?
Reference Passages: [1] reporting that the Los Angeles-based company is aiming to raise \$500 million at a valuation of \$16 billion to \$19 billion, making it the third most highly valued tech start-up backed by venture capitalists. The Chinese handset maker Xiaomi is valued at \$45 billion, while Uber is estimated to be valued at about \$40 billion, according to data from CB Insights. Read MoreVC investment hits \$86B thanks to Uber, Xiaomi Snapchat was valued at \$10 billion in August, according to a Dow Jones report. Some of its investors from previous rounds include Benchmark, Lightspeed Venture Partners and Kleiner Perkins Caufield
[2] Ever wondered how highly valued technology giants like Google and Facebook were able to grow so fast and pay their employees so well in such a short amount of time, or how still growing start-ups like Uber are able to lose 1.2 billion US dollars in just the first half of this year alone and still command a valuation upwards of 50 billion US dollars? The answer lies with a special category of investment activity known as venture capital. Venture capitalists are professional investors who invest in a number of highly scalable high-risk technology ventures hoping to make a multi-fold
[3] …
Answer: Yes. Did you watch The Social Network? They went a while before introducing ads, so they could make money, as they needed to establish their brand and amass users. Once you have dedicated users, introducing ads won’t deter most, but if you are still new, having ads will deter a lot. The same goes for Uber, it’s not that they aren’t making money, it’s that they are reinvesting a ton of it to make their service better.
Table 9: Example of the test datasets.
C.2 Details of Methods
Prompting
[…]
ReClaim w/IG
@@ C.2 Details of Methods
1 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
@@
3 ReferModel Instrction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …

### v2

@@ Methods
EM Rec. Exact match recall rate of gold short answers in the text generated by the LLM.
Claim Rec. Recall rate of generated claims in the text generated by the LLM.
@@
Table 5: The notation table.
In Table 5, we list the notations and abbreviations in this paper, together with their definitions.
@@ Appendix B Details of Training Dataset Construction
B.1 Details of Training Dataset
The statistics of the training dataset is shown in Table 1
WebGLM-QA
The WebGLM-QA Liu et al. (2023) dataset is a pioneering resource designed to bolster the development and assessment of web-enhanced question answering systems. It distinguishes itself by seamlessly integrating web search functionalities into the QA process, empowering systems to tap into the expansive repository of knowledge on the internet. Carefully curated to overcome the shortcomings of existing datasets, WebGLM-QA offers a holistic and pragmatic solution for open-domain question answering tasks. It consists of 43,579 high-quality data samples for the train split, 1,000 for the validation split, and 400 for the test split. Refer to our paper for the data construction details.
In WebGLM-QA, the provided answers is generated by the model based on the given reference passages; therefore, we directly use it as our gold answers.
ELI5
ELI5 Fan et al. (2019) dataset is a benchmark in natural language processing designed for long-form question answering tasks, focusing on complex and explanatory questions (see Table LABEL:table:testdatacase). It comprises 270K threads from the Reddit forum “Explain Like I’m Five” (ELI5) where an online community provides answers to questions which are comprehensible by five year olds. For ALCE Gao et al. (2023b), ELI5 questions were paired with passages from Sphere Piktus et al. (2021), a filtered version of Common Crawl.
For the ELI5 dataset, due to the distributional differences between human answers and the retrieved reference passages, we opted to use Llama-3.1-405B-Instruct to generate answers based on the top-5 reference passages as the gold answers.
B.2 Citation Selection
During the training data structuring phase, we initially employ Llama-3.1-405B-Instruct to automatically segment the answers within the WebGLM-QA dataset and ELI5 dataset into clauses. Subsequently, for each of these clauses, we undertake a search for corresponding sentence-level, fine-grained citations from the provided reference passages. The prompt we use is as follows:
@@ B.3 Citation Filtering
Appendix C Experiment Settings
C.1 Details of Datasets
The statistic of the test datasets is shown in Table 9. In Figure 5 and Figure 6, we employ bge-v2-m3 Li et al. (2023); Chen et al. (2024) to compute the relevance between queries and reference passages, presenting the statistical information of both training and test dataset in the form of box plots. Below, we provide a detailed description of the test datasets.
ASQA
ASQA Stelmakh et al. (2022) is the first long-form question answering dataset that focuses on ambiguous factoid questions (see Table LABEL:table:testdatacase). It contains 4,353 samples for the train split and 948 samples for the dev split. For ALCE Gao et al. (2023b), ASQA questions were paired with passages from Wikipedia passages (2018-12-20 snapshot) which purportedly contained the answers.
EXPERTQA
EXPERTQA Malaviya et al. (2024) dataset is a benchmark designed for evaluating the factuality and attribution capabilities of large language models across diverse domains. It contains 2177 long-form questions curated by experts from 32 fields, along with LLM-generated answers that have been revised by these experts to ensure factuality and proper sourcing. The dataset aims to provide a high-quality resource for developing and assessing AI systems that can deliver accurate and well-referenced information tailored to the needs of domain specialists.
ELI5
Descriptions can be found in the section B.1
C.2 Details of Methods
Prompting
ReClaim ${}_{\texttt{Unified}}$
In ReClaim ${}_{\texttt{Unified}}$ setting, we directly train the model using the constructed training dataset in Appendix B and employ the same instruction as in the prompting setting for one-step generation of attributed text.
ReClaim w/IG
C.3 Details of Evaluation Metrics
EM Rec.
This metric is employed to assess the accuracy of the ASQA dataset. It utilizes a list of short answers (typically in the form of words or phrases) provided within the ASQA dataset to perform string matching within the generated long answers generated. The proportion of short answers that can be matched relative to the total length of the short answer list is calculated to determine the metric.
Claim Rec.
This metric is designed to evaluate the accuracy of the ELI5 and EXPERTQA datasets. Since these datasets do not provide short answer lists but only a long-form answer, ALCE employs InstructGPT Ouyang et al. (2022) to sample three sub-claims from the standard answer. Then use the NLI model to assess the entailment relationship between each sub-claim and the LLM-generated response. The final score for this metric is determined by the proportion of sub-claims that are entailed in the LLM’s answer.
MAUVE
We use MAUVE (Measuring the gap between neural text and human text using divergence frontiers) to measure the fluency of the generated text. MAUVE is a statistical metric that quantifies the similarity between neural-generated text and human-written text by computing the divergence frontiers between them. A higher MAUVE score indicates that the generated text is more coherent and natural, closely resembling human language style.
CAS
This metric, consistent with the Citation Recall metric in ALCE, is designed to evaluate the proportion of answer clauses that are sufficiently supported by citations. Specifically, for each answer clause associated with a citation, a Natural Language Inference (NLI) model is employed to determine whether an entailment relationship exists between the clause and its corresponding citation. If such a relationship is identified, it indicates that the citation adequately substantiates the answer clause. The final score for this metric is derived from the ratio of answer clauses that exhibit an entailment relationship with their citations.
CRS
This metric aligns with the Citation Precision metric in ALCE, designed to assess whether the citations contain redundancy. In the ALCE, citations are identified by paragraph numbers, and redundancy is evaluated by sequentially removing paragraphs. However, in our approach, to accommodate sentence-level citations, we first segment the citation into sentences using NLTK. We then iteratively remove sentences and verify whether the remaining text still maintains an entailment relationship with the answer clause. Finally, the proportion of non-redundant citations is used as the CRS score.
C.4 Training Details
We trained the model using llama-factory Zheng et al. (2024) on four A800 80G GPUs. For the Unified approach, we applied full fine-tuning with a learning rate of 3e-5, 3 epochs, and a total batch size of 128. For the ReClaim w/IG approach, we used LoRA to train two models: ReferModel with a learning rate of 5e-5, 5 epochs, and a total batch size of 128, and ClaimModel with a learning rate of 3e-5, 2 epochs, and a total batch size of 128. During model inference, we can alternate between two LoRA parameters to reduce memory requirements.
Figure 5: Reranker score distribution between query and reference passages in the training dataset.
Figure 6: Reranker score distribution between query and reference passages in the test datasets.
Appendix D Additional Results
D.1 Default Datasets
For our test datasets, in addition to providing gold passages or reranked passages as reference passages, we also tested the results of directly using the top-5 retrieved passages, as shown in Table 6.
Dataset Type Fluency Correct. CAS CRS
ASQA Oracle 88.1 53.5 92.1 71.6
Default 88.1 40.0 91.0 84.4
ELI5 Oracle 71.6 17.8 89.9 67.5
Default 75.7 6.5 89.4 66.3
EXPERTQA Rerank 63.5 14.0 90.1 68.6
Default 62.9 12.2 87.2 65.7
Table 6: Results of the default test datasets.
The results reveal that although the accuracy of the LLM’s responses decreases due to varying quality of reference passages provided to the LLMs, it still maintains high CAS and CRS scores. This indicates that our method maintains strong attribution performance even with passages retrieved directly.
D.2 Additional Baselines
Due to the differences in datasets and evaluation metrics, we include this baseline in the Appendix.
Learning to Plan and Generate Text with Citations Fierro et al. (2024) conceptualizes plans as a sequence of questions that serve as blueprints for content generation and organization. Two variants of blueprint models are introduced: an abstractive model, where questions are generated from scratch, and an extractive model, where questions are copied from input passages.
For the comparative analysis of this method, we utilized the ASQA and ELI5 datasets, with top-5 reference passages retrieved directly via BM25 and GTR. The experimental results are shown in Table 7.
ASQA ELI5 Average
C A C A C A
LongT5 3B (10-psg) +$Blueprint_{A}$ +Attribution 33.8 77.8 5.2 60.9 19.5 69.4
 ReClaim w/IG (5-psg) 40.0 87.6 6.5 76.1 23.3 81.9
 ReClaim w/IG (10-psg) 40.8 88.5 8.0 72.9 24.4 80.7
Table 7: Comparison with the baseline: Learning to Plan and Generate Text with Citations. C denotes accuracy, while A represents the attribution score, which is the F1 score of CAS and CRS.
The experimental results indicate that our approach outperforms this method in terms of both answer accuracy and attribution scores.
D.3 Faithfulness Results
The experimental results of the faithfulness metric for the ASQA and EXPERTQA datasets are shown in Figure 7 and Figure 8.
Figure 7: The faithfulness analysis results of the ASQA dataset.
Figure 8: The faithfulness analysis results of the EXPERTQA dataset.
Figure 9: The comparative experimental results of the ASQA dataset.
Figure 10: The comparative experimental results of the EXPERTQA dataset.
@@ D.3 Faithfulness Results
ASQA 948 Factoid (ambiguous) Wikipedia (21M) 9.0 517.5 71.8
ELi5 1000 Why/How/What Sphere (899M) 16.5 546.3 121.5
EXPERTQA 1000 ambiguous/unambiguous Sphere (899M) 19.4 600.0 152.2
@@
Table 11: Example of the test data.
Question: When data is compressed or zipped, what is actually happening to the data?
Reference Passages:
[1] Title: Lossless vs. Lossy Compression: What’s the Difference?
Text: data in several different ways, balancing fidelity and efficiency for functional and presentable data on the egress end. A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources. The model is able to compress and decompress digital media such that the output perfectly matches the input. The Zip file archival tool is a well-known format that supports lossless compression. Zip files compress digital media into much smaller data (often given the 2̆018.zip2̆019 file extension) that can be uncompressed into their original form,
[2] Title: What is MP3? (Designing Web Audio)
Text: space, try encoding in mono. To make sure you trap all possible spatial data, encode in stereo mode. Most users find that joint stereo is adequate for most purposes. As mentioned earlier, compressing a WAV file with zip doesn’t shave much off the file size, which is why psychoacoustics are employed. However, the MP3 encoding process actually does employ the classic Huffman encoding algorithm. After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern. It’s as though zip-type encoding were being run internally on the psychoacoustically encoded data
[3] Title: US Treasury: addendum | Zip (File Format) | Comma Separated Values
Text: most large files on the Internet are compressed. File compression reduces the size of a file and the time it takes to download. Compression software uses complex mathematical equations to scan a file for repeating patterns in the data. It replaces the data with smaller codes that take up less room. This can dramatically decrease the size of report files. Viewing a 2̆01cZipped2̆01d file To view compressed files, you need a compatible decompression program that can read these codes and convert the data back to its original form. To decompress a .zip file you need a utility like WinZIP2̆122 (an
[4] Title: An introduction to data compression
Text: data and try to find redundant parts of that data and remove those redundancies. That wasn’t too simple, but the definition needs to be somewhat complex in order to differentiate compression from some other technologies. 2. How is compression accomplished? Can you give us a picture of what actually happens to the data? Basically, they’re trying to find common pieces of data blocks that they can get rid of, shrink, remove or substitute with smaller patterns. The more of those things they can find, the more it can compress. A perfect example to illustrate how that works is if you
[5] Title: Introduction to Data Compression - Data Compression | Coursera
Text: we get into the topic. so, just specific applications that, maybe are familiar for file compression. All file systems and, and disks have built-in, compression technologies. Such as, as zip, and b-zip and many others of similar type. Technologies. And the multimedia, everybody’s familiar with, the JPEG and MP3 and MPEG, and all those sorts of things for images, sound and video. Those are all about data compression. and for communication, now, that’s, what has, what enabled, fax, and also enables new technologies, like Skype, the ability to, reduce the, amount of data that you actually need to send. for (Irrelevant passage)
Gold Answer: For most algorithms, the compressor will search for a pattern that appears multiple times within the data, and then replace that pattern with a shorter symbol, and create an entry in a "dictionary" recording what each symbol means. Compressing algorithms can get a little more complex than that, but that’s the general idea.
Model Answer: According to the citation: <reference> Compression software uses complex mathematical equations to scan a file for repeating patterns in the data. It replaces the data with smaller codes that take up less room. </reference> We can know that: <claim> Compression software uses complex mathematical equations to scan a file for repeating patterns in the data and replaces the data with smaller codes that take up less room. </claim> According to the citation: <reference> A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources. </reference> We can know that: <claim> A common implementation of lossless file-compression includes the use of Huffman coding, whose redundancy-limiting algorithm recognizes patterns in groups in order to conserve time, space and other resources. </claim> According to the citation: <reference> After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern. </reference> We can know that: <claim> After all psychoacoustic methods have been applied, the Huffman encoding pass seeks out and compresses any remaining redundancies in the bit pattern. </claim> According to the citation: <reference> Basically, they’re trying to find common pieces of data blocks that they can get rid of, shrink, remove or substitute with smaller patterns. The more of those things they can find, the more it can compress. </reference> We can know that: <claim> The more of these common pieces of data blocks that can be found, the more it can compress. </claim>
@@
1 ReferModel Instruction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …
@@
3 ReferModel Instruction: Given the Question and References below, provide an answer for the Question that is generated using information exclusively from the References(some may be irrelevant). …